Eliasbeth is 6 years older than her sister elisa. Their mother's age is twice the sum of their ages. How old are they if their mother is 40?
Answer by sudhanshu_kmr(1152)


Elisa age = x
Eliasbeth age = x + 6
mother's age = 2( x + x+6)
40 = 2( 2x +6)
=> 40 = 4x + 12
=> 4x = 40-12 = 28
=> x = 7


Elisa age = 7, elisabeth age = 13